OpenStudy (anonymous):

The graph below shows the revenue earned by a company in its first four quarters of operation. A line of best fit has been drawn on the graph.

OpenStudy (anonymous):

Based on the linear model, how much revenue, in thousands of dollars, should the company earn in its 7th quarter of operation? Type your answer in the box below.

OpenStudy (wolf1728):

Hello

OpenStudy (anonymous):

Hi :-)

OpenStudy (wolf1728):

Hard to judge from that graph but I'd say about 70,000

OpenStudy (anonymous):

Okay this one kind of confused me too

OpenStudy (anonymous):

But I'll be right back I will get at you if I need help again thanks so much

OpenStudy (anonymous):

Lagging so I cannot even see what you wrote @doc.brown

OpenStudy (wolf1728):

Gee I answered that as $70,000 (I just continued the line on the graph). I think Doc Brown's answer stayed posted fo about two seconds.

OpenStudy (anonymous):

Oh lol btw you are right wolf

OpenStudy (doc.brown):

I forgot to carry the something or other and got the wrong answer.

OpenStudy (wolf1728):

thanks hellogoodmorning

OpenStudy (anonymous):

No worries thanks for helping and No problem Wolf

OpenStudy (doc.brown):

\[\frac{\Delta x}{\Delta y}\]\[\frac{4-1}{39-11}\]\[\frac{3}{28}\]\[7\times\frac{28}{3}\]\[=65\]thousand dollars

OpenStudy (wolf1728):

Well almost 6:00 a.m. here Think I'll get some sleep. Good night folks. (or maybe that's more like good morning?)

OpenStudy (anonymous):

Loll Yes HelloGoodnight to you too

OpenStudy (anonymous):

and @doc.brown you think it's 65k